### v1

# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations via Projected Gradient Descent (PGD), and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and VQA tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks.
1 Introduction
Multimodal Large Language Models (MLLMs) have made significant strides in integrating vision and language understanding within a unified architecture  llava ; instructblip ; qwen25vl . By combining powerful visual encoders and large language models through alignment mechanisms, MLLMs such as LLaVA llava and Qwen-VL qwen25vl demonstrate strong capabilities across a wide range of multimodal tasks. However, beneath their seemingly impressive performance lies a critical limitation: MLLMs often fail to distinguish between relevant and irrelevant signals across modalities, leading to unreliable predictions wang2024mdpo ; vlmconflict ; spurlens . Moreover, while MLLMs are designed for multimodal tasks, their failure on unimodal tasks—where only a single modality (e.g. text) should guide the prediction—raises concerns about whether the model can preserve modality-specific competencies. For instance, MLLMs frequently underperform on pure visual recognition imagewikiqa ; mof and textual reasoning vlmconflict ; cogvlm ; vila , suggesting that cross-modal fusion may induce unintended interference and degrade unimodal performance.
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting imagewikiqa ; mof ; cogvlm ; vila , knowledge conflict vlmconflict ; wang2024mdpo , and spurious correlations more ; spurlens . Catastrophic forgetting has been identified as a key factor in visual degradation imagewikiqa ; mof , where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict vlmconflict ; wang2024mdpo impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations more ; spurlens show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA llava fail to fully align vision and language representations, resulting in unstable modality reliance mof ; vlmconflict . Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding imagewikiqa and diminished language understanding vila . Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency Competency —the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
Figure 1: Perturbation results on image classification (left) and text QA (right) tasks. Left: Mini-ImageNet with Orig (Original input), UF (Add Unrelated Facts), and MD (Add Misleading Descriptions). Right: OpenBookQA with RP (Random Pixels), IR (Attach Irrelevant Real Image), BK (Attach Black Canvas), and WH (Attach White Canvas). For image-heavy tasks, spurious textual cues (esp. misleading descriptions) severely degrade accuracy. For text-heavy tasks, unrelated real images disrupt reasoning—highlighting modality interference in MLLMs.
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles more to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure 1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images significantly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
The empirical results from Figure 1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentations strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations includes both heuristic perturbations (e.g. add unrelated facts in image-heavy prompts) and adversarial training time perturbations, which exposes the model to worst-case alignment disruptions, serving as a strong variant. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families llava15 ; instructblip ; qwen25vl and diverse benchmarks miniimagenet ; caltech101 ; openbookqa ; mmlu ; scienceqa ; mmbench ; seedbench demonstrate the superiority of our method.
2 Related Works
Improving Modality Alignment in Multimodal Language Models
Recent studies have revealed that modality misalignment remains a key obstacle in MLLMs, leading to degraded performance on both image-heavy and text-heavy tasks. For visual understanding, catastrophic forgetting occurs when multimodal tuning overrides pretrained visual features imagewikiqa ; mof ; cogvlm . In text-heavy scenarios, knowledge conflict vlmconflict arises when inconsistent parametric knowledge from different modalities confuse reasoning. Moreover, mDPO wang2024mdpo identifies language bias in training, where models fail to condition their responses on visual input. Some works attribute such issues to shallow fusion cogvlm —e.g., LLaVA llava uses lightweight projectors to bridge vision and language spaces, leaving a representational gap and resulting in loosely coupled features mof ; vlmconflict . Others highlight data limitations: even well-encoded visual features fail to support reasoning without adequate supervision to guide decoding imagewikiqa .
Building on these diagnoses, recent models have proposed multiple solutions. MoF mof mitigates this by fusing features from multiple vision encoders, while VLMClassifier imagewikiqa enhances recognition via vision-only finetuning, though it struggles with VQA due to lack of cross-modal alignment. CogVLM cogvlm introduces a visual expert module to improve vision-language integration. VILA vila , QwenVL qwen25vl , and InternVL internvl3 incorporate text-only supervision in different ways to preserve or enhance language capabilities during multimodal training—through stage-wise separation, parallel preservation, and unified joint optimization, respectively. These works motivate us to hypothesize on the root cause—the model’s inability to assess modality relevance. We further propose a causal framing of modality interference and introduce a perturbation-based fine-tuning strategy to improve modality discriminability and inference-time robustness.
Adversarial Robustness Across Modalities
Adversarial perturbations threaten the reliability of both vision and text tasks by exposing vulnerabilities through deliberate and imperceptible perturbations. Existing robustness methods can be broadly categorized by modality, targeting either continuous image embeddings or discrete token spaces.
In vision tasks, attacks like FGSM DBLP:journals/corr/GoodfellowSS14 and CW DBLP:conf/sp/Carlini017 first revealed the fragility of neural networks to imperceptible input changes. PGD DBLP:conf/iclr/MadryMSTV18 formalized this under a saddle-point framework, becoming the standard for adversarial training. AutoAttack DBLP:conf/icml/Croce020a further unified strong attacks, including PGD variants, into a reliable benchmark. In text tasks, adversarial methods must contend with discrete inputs. TextFooler DBLP:conf/aaai/JinJZS20 substitutes key words with semantically similar ones to mislead predictions, while CodeAttack DBLP:conf/aaai/JhaR23 adapts this idea to code-language models. More recently, PGD has been extended to LLMs via continuous relaxation DBLP:journals/corr/abs-2402-09154 , enabling efficient, gradient-based attacks in embedding space. Beyond evaluation, PGD has also been used as a regularizer to improve optimization. PTP DBLP:conf/emnlp/ChenCHC23 applies PGD-style perturbations in the prompt embedding space to smooth training and enhance stability. Inspired by this, our work extends PGD to the multimodal embedding space, enabling unified gradient-based control over both visual and textual inputs.
3 Causal Analysis on Modality Interference
Cross-Modality Competency Problems in Multimodal Large Language Models
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features Competency . We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency Problem as a class of tasks where no single modality should be sufficient to determine the correct prediction, and models are expected to integrate information across modalities to make robust decisions. A competent model should be able to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
Figure 2: Causal graph illustrating modality interference in our perturbation-based evaluation analysis. Controlled interventions (heuristic) perturb either the image or text inputs, affecting their intermediate representations and ultimately the model prediction.
Perturbation-based Evaluation Experiment
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section 5 and Section C.1.
Causal Modeling of Modality Interference
We formalize modality interference through a causal intervention perspective with a causal graph, as shown in Figure 2, where visual inputs ($X_{I}$) and textual inputs ($X_{T}$) are processed into their respective representations ($Z_{I}$, $Z_{T}$) before being fused to produce the final prediction ($A$). To study the model’s reliance on different modalities, we introduce perturbations directly at the input level, serving as causal interventions causalIntervention on $X_{I}$ and $X_{T}$. Under ideal cross-modality competency, the model’s prediction should primarily depend on the task-relevant pathway (e.g., $X_{I}\rightarrow Z_{I}\rightarrow A$ in image-heavy tasks, $X_{T}\rightarrow Z_{T}\rightarrow A$ in text-heavy tasks). Causal interventions at the input level allow us to diagnose whether the model improperly fuses irrelevant signals into its decision process. We use $x_{I}\textquoteright$ ’ to denote an intervention on image $X_{I}$ and use $x_{T}\textquoteright$ ’ as an intervention on text $X_{T}$. Following Pearl’s causal framework causalIntervention ; intermediaryVariable , we quantify the impact of modality perturbations on model predictions by formalizing causal effects in our multimodal setting. Specifically, we define the pre-intervention prediction distribution as $P(A|X_{I},X_{T})$, and the post-intervention prediction distribution after applying a perturbation on $X_{I}$ or $X_{T}$ as $P\textquoteright(A|\text{do}(X_{I}=x_{I}\textquoteright)$ ’ ) or $\text{do}(X_{T}=x_{T}\textquoteright))$ ’ ) ). The do-operation represents an intervention to specific modality, and the causal effect (CE) of an intervention is evaluated via a distance metric $\delta$ comparing $P$ and $P\textquoteright$ ’ as $\text{CE}=\delta(P,P^{\prime})$. We assess the causal effect via prediction changes using $\delta_{\text{cp}}(P,P^{\prime}):=\mathbb{I}\left(a\neq a^{\prime}\right)$ in which $a=\arg\max_{x}P(x)$ is the predicted answer before intervention, $a\textquoteright=\arg\max_{x}P\textquoteright(x)$ is the predicted answer after intervention and $\mathbb{I}(\cdot)$ is the indicator function that outputs $1$ if $a\neq a\textquoteright$ ’ and 00 otherwise. Thus, $\delta_{\text{cp}}$ captures whether the model’s final decision $A$ changes under perturbations to the input modality. In all interventions, a high value of $\delta_{\text{cp}}$ indicates the model’s susceptibility to modality interference, revealing spurious reliance on irrelevant modality.
4 Methods
To mitigate modality interference and enhance cross-modality competency, we propose a unified perturbation-aware training framework that introduces interventions at both the input level (on $X_{I}$ and $X_{T}$) and the representation level (on $Z_{I}$ and $Z_{T}$) with consistency regularization. Overall pipeline is displayed in  Figure 3.
Figure 3: Overview of our proposed framework.
4.1 Perturbation-based Data Augmentation
To increase causal robustness along the desired paths $X_{I}\rightarrow Z_{I}\rightarrow A$ (for image-heavy tasks) or $X_{T}\rightarrow Z_{T}\rightarrow A$ (for text-heavy tasks), we first adopt a diverse and causally grounded data augmentation method by augmenting each sample with both heuristic perturbations and training time adversarial perturbations.
Mixture of Multi-Task Training data with Heuristic Perturbations
Let $\mathcal{D}={(x_{I},x_{T},a)}$ denote the full multimodal training dataset, where $(x_{I},x_{T})$ are the image and text inputs and $a$ is the ground-truth answer. The dataset $\mathcal{D}$ can be partitioned into three subsets based on tasks: 1) Image-heavy set $\mathcal{D}^{\text{img}}$: samples where visual input $x_{I}$ is the dominant information source; 2) Text-heavy set $\mathcal{D}^{\text{text}}$: samples where textual input $x_{T}$ provides the main reasoning signal; 3) VQA set $\mathcal{D}^{\text{vqa}}$: samples from vision-language datasets with naturally balanced multimodal dependencies. In practice, we transform these image-heavy and text-heavy datasets into VQA format to construct $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$, and derive VQA samples from the supervised finetuning stage of each MLLM as $\mathcal{D}^{\text{vqa}}$. For each sample $(x_{I},x_{T},a)\in\mathcal{D}^{\text{img}}\cup\mathcal{D}^{\text{text}}$, we maintain its original version and apply heuristic perturbation to construct Origin Samples and Perturbation-Augmented Samples. Origin samples are used to reinforce the desired causal path (e.g., $X_{I}\rightarrow Z_{I}\rightarrow A$). Perturbation-augmented samples are variants of the same instance with perturbations applied to the irrelevant modality, which are denoted as $(x_{I},\tilde{x}_{T},a)\in\mathcal{D}^{\text{img}}_{\text{pert}}$ and $(\tilde{x}_{I},x_{T},a)\in\mathcal{D}^{\text{text}}_{\text{pert}}$ where $\tilde{x}_{T}$ and $\tilde{x}_{I}$ are perturbed text and image respectively. Together, the augmented dataset can be written as:
$\mathcal{D}^{\text{AUG}}=\mathcal{D}^{\text{img}}\cup\mathcal{D}^{\text{img}}_{\text{pert}}\cup\mathcal{D}^{\text{text}}\cup\mathcal{D}^{\text{text}}_{\text{pert}}\cup\mathcal{D}^{\text{VQA}}.$ (1)
We sample $N_{\text{img}}$ and $N_{\text{text}}$ examples from $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$ to construct $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$ respectively, and the remaining $N_{\text{vqa}}$ examples are VQA samples from $\mathcal{D}^{\text{vqa}}$ to construct $\mathcal{B}^{\text{vqa}}$ With dynamically generated perturbed variants for each image-heavy and text-heavy sample, the final training batch is:
$\mathcal{B}=\mathcal{B}^{\text{img}}_{\text{orig}}\cup\mathcal{B}^{\text{img}}_{\text{pert}}\cup\mathcal{B}^{\text{text}}_{\text{orig}}\cup\mathcal{B}^{\text{text}}_{\text{pert}}\cup\mathcal{B}^{\text{vqa}},$ (2)
where $\mathcal{B}^{\text{img}}_{\text{pert}}$ and $\mathcal{B}^{\text{text}}_{\text{pert}}$ are perturbation-augmented variants generated from $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$, respectively. For the full training batch $\mathcal{B}$, which includes both original and perturbed samples, we define the supervised loss $\mathcal{L}_{\text{sft}}$ as the cross-entropy loss computed over all answer tokens in the ground-truth sequences. Let $\mathcal{L}_{\text{cls}}(x_{I},x_{T},a)$ denote the standard autoregressive loss for a sample $(x_{I},x_{T},a)$, then:
$\mathcal{L}_{\text{sft}}=\frac{1}{|\mathcal{B}|}\sum_{(x_{I},x_{T},a)\in\mathcal{B}}\mathcal{L}_{\text{cls}}(x_{I},x_{T},a).$ (3)
This objective enables full supervised fine-tuning across modality-specific samples with their heuristic perturbed variants and VQA samples, encouraging the model to jointly learn robust unimodal reasoning and generalizable multimodal alignment.
Adversarial Perturbation with Cross-modality Masking
While heuristic perturbations simulate realistic but limited modality noise at the input level, they may not fully capture the worst-case failure modes of MLLMs, especially under complex spurious alignments in the representation space. To overcome this limitation, we introduce a stronger and more generalizable intervention through adversarial training. These perturbations simulate worst-case alignment disruptions during training, serving as targeted interventions on latent nodes $(Z_{I},Z_{T})$ to reveal the Direct Causal Effect (DCE) of irrelevant modalities on $A$. By optimizing the model under such adversarial conditions, we reduce the model’s reliance on spurious cross-modal signals and reinforce task-relevant causal pathways. Specifically, we apply Projected Gradient Descent (PGD) DBLP:conf/iclr/MadryMSTV18 ; ptp at training time to perturb the multimodal token embeddings $(Z_{I},Z_{T})$ directly. It constructs perturbations $\delta=(\delta_{I},\delta_{T})$ in the latent space that maximize the model’s predictive loss:
$\delta=\underset{\|\delta\|_{\infty}\leq\epsilon}{\text{argmax}}\,\mathcal{L_{\text{cls}}}(f(Z_{I}+\delta_{I},Z_{T}+\delta_{T})),$ (4)
where $\epsilon$ bounds the perturbation strength and $f$ denotes the model’s prediction function. We implement this objective via $n$-step iterations, where at each step $t$, the perturbation is updated as:
$\delta^{(t+1)}=\Pi_{\|\delta\|_{\infty}\leq\epsilon}\left(\delta^{(t)}+\alpha\cdot\left(\nabla_{\delta}\mathcal{L_{\text{cls}}}(f(Z+\delta^{(t)}))\right)\right),$ (5)
where $\Pi$ is the projection operator onto the $\ell_{\infty}$-ball, $\alpha$ is the step size, and $Z=[Z_{I};Z_{T}]$ is the concatenated embedding. By injecting these adversarial perturbations into $(Z_{I},Z_{T})$ during training, we simulate worst-case interventions on the model’s internal decision pathways, exposing and correcting spurious dependencies between irrelevant modality features and prediction $A$. In practice, we restrict the perturbation to the task-irrelevant modality using a modality-specific binary mask $M\in\{0,1\}^{L\times d}$, where $L$ is the sequence length and $d$ is the dimension. Let $\mathbf{E}\in\mathbb{R}^{L\times d}$ denote the multimodal token embedding sequence, where projected visual features are already interleaved into the textual token sequence llava ; qwen25vl ; internvl3 . We construct adversarially perturbed embeddings $\tilde{\mathbf{E}}$ as:
$\tilde{\mathbf{E}}=\mathbf{E}+\delta\odot M,$ (6)
where $\delta\in\mathbb{R}^{L\times d}$ is the perturbation and $\odot$ denotes element-wise masking. We then optimize $\delta$ via $T$ steps of projected gradient ascent:
$\delta^{(t+1)}=\Pi_{\|\delta\|_{\infty}\leq\epsilon}\left(\delta^{(t)}+\alpha\cdot\left(\nabla_{\delta}\mathcal{L}(f(\mathbf{E}+\delta^{(t)}\odot M))\right)\right),$ (7)
where $\alpha$ is the step size and $\Pi$ projects the noise back into the $\ell_{\infty}$ ball with radius $\epsilon$, we initialize $\delta$ with Gaussian noise $\mathcal{N}(0,\epsilon^{2})$. After obtaining $\tilde{\mathbf{E}}$, we compute the adversarial loss as $\mathcal{L}_{\text{adv}}=\mathcal{L_{\text{cls}}}(f(\tilde{\mathbf{E}}))$.
4.2 Consistency Regularization under Perturbations
While perturbation-based data augmentation exposes the model to diverse interventions, it does not explicitly constrain how the model’s intermediate representations ($Z_{I}$, $Z_{T}$) should respond to such perturbations. In multimodal settings, even small interventions on the task-irrelevant modality can undesirably shift the fused representations and lead to inconsistent predictions. To address this, we introduce a consistency regularization strategy that enforces output stability between original and perturbed inputs, serving as an indirect but effective constraint on the representations $Z_{I}$ and $Z_{T}$ to mitigate modality interference. By minimizing the divergence between the prediction distributions of original and perturbed inputs, the model is encouraged to maintain invariant behavior along the task-relevant causal paths. Specifically, to illustrate the role of consistency regularization, consider an image-heavy sample $(x_{I},x_{T},a)\in\mathcal{B}^{\text{img}}_{\text{orig}}$ and its perturbed version $(x_{I},\tilde{x}_{T},a)\in\mathcal{B}^{\text{img}}_{\text{pert}}$. The model’s predictive distribution over the answer $a$ is denoted as $p_{\theta}(a|x_{I},x_{T})$, parameterized by $\theta$. We define the consistency loss between the original and perturbed inputs as:
$\mathcal{L}_{\text{consistency}}=\mathrm{Consistency}\left(p_{\theta}(A|x_{I},x_{T})\,\|\,p_{\theta}(A|x_{I},\tilde{x}_{T})\right)$ (8)
This objective encourages the model to rely primarily on the visual representation $Z_{I}$ and to be invariant to perturbations in $X_{T}$, thus regularizing the causal path $Z_{I}\rightarrow A$ against interference. In practice, we calculate adopt multiple methods (e.g. KL, JS) based on the distribution of the output logit of the model produced by its language model head. Let $l^{\text{orig}}\in\mathbb{R}^{L_{A}\times V}$ and $l^{\text{pert}}\in\mathbb{R}^{L_{A}\times V}$ denote the pre-softmax logits produced by the language model for the original and perturbed samples respectively, where $L_{A}$ is the number of answer tokens and $V$ is the vocabulary size. We define the token-level consistency loss with temperature $\tau$ as (using KL divergence for illustration):
$\mathcal{L}_{\text{consistency}}=\frac{1}{L_{A}}\sum_{i=1}^{L_{A}}\sum_{v=1}^{V}\text{softmax}\left(\frac{l^{\text{orig}}_{i}}{\tau}\right)_{v}\cdot\log\left(\frac{\text{softmax}\left(\frac{l^{\text{orig}}_{i}}{\tau}\right)_{v}}{\text{softmax}\left(\frac{l^{\text{pert}}_{i}}{\tau}\right)_{v}}\right).$ (9)
The same principle applies in text-heavy tasks, where perturbations are instead applied to the image input $X_{I}$, and consistency is enforced between $p_{\theta}(A\mid x_{I},x_{T})$ and $p_{\theta}(A\mid\tilde{x}_{I},x_{T})$.
4.3 Final Training Objective
Our final training objective integrates both perturbation-based data augmentation and consistency regularization into a unified framework. For each batch, we begin with a set of original samples $\mathcal{B}_{\text{orig}}$ and dynamically construct their heuristic perturbed counterparts $\mathcal{B}_{\text{pert}}$ via input-level augmentations. We then apply PGD-based adversarial perturbations on both $\mathcal{B}_{\text{orig}}$ and $\mathcal{B}_{\text{pert}}$, and enforce consistency between the predictions of original and all perturbed samples. Specifically, we define the overall loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sft}}+\mathcal{L}_{\text{adv}}+\lambda_{\text{cons}}\cdot\mathcal{L}_{\text{consistency}},$ (10)
where $\mathcal{L}_{\text{sft}}$ is the supervised loss computed over $\mathcal{B}$, $\mathcal{L}_{\text{adv}}$ is the adversarial loss computed on all adversarial perturbed samples and $\mathcal{L}_{\text{consistency}}$ is the consistency loss between original and all perturbed sample pairs. By aligning all three losses with the causal structure of multimodal reasoning, we systematically mitigate modality interference and improve cross-modality competency in MLLMs.
5 Experiments
Models. We conduct experiments on three MLLM familiies with different parameter size: Qwen2.5-vl-3b qwen25vl , LLaVA-1.5-7B & LLaVA-1.5-13B llava15 and InstructBLIP-Vicuna-7B instructblip . These models typically integrate a pre-trained vision encoder (e.g., CLIP ViT-L/14 clip ) with a language model backbone (e.g., Vicuna vicuna ), connected via a modality projection module. Following the original LLaVA and Qwen-VL implementation, we freeze the vision encoder and train the multimodal projector and the language model. In our adaptation of InstructBLIP, we adopt an alternative tuning strategy: we freeze both the vision encoder and the Q-Former, and fine-tune the language model and the linear projection layer. We provide further discussion in  Appendix B.
Baselines. We include following baselines for comparison: LLaVA-1.5-13B + I-MoF mof : By applying the designed Interleaved Mixture-of-Features (I-MoF) module on LLaVA-1.5-13B llava15 to spatially combine CLIP clip and DINOv2 dinov2 visual tokens, it enhances visual grounding by integrating complementary features from contrastive and self-supervised vision encoders. VLMClassifier imagewikiqa : it enhances visually-grounded language models for image classification by fine-tuning them on ImageNet imagenet (VLMClassifier-1) or ImageNet combining LLaVA-Instruct llava (VLMClassifier-2).
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet miniimagenet and Caltech-101 caltech101 , used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA openbookqa and MMLU mmlu , consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset llava as the instruction-tuning dataset for related models. For InstructBLIP, we additionally include TextCaps textcaps , since these are the only publicly accessible VQA datasets known to be used in instruction-tuning for related models. For Qwen2.5-VL, whose instruction-tuning data is in-house and not publicly available, we use LLaVA-Instruct-dataset as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG scienceqa , MM-Bench-EN mmbench , and Seed-Bench-IMG seedbench . For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in the  Appendix D.
Table 1: Evaluation of unimodal and multimodal tasks across multiple methods. For unimodal datasets, we report accuracy on origin setting (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on the original setting. The best accuracy is marked in bold.
Model Settings Mini-ImageNet Caltech-101 OpenBookQA MMLU ScienceQA MM-Bench Seed-Bench
Orig Perturbed Orig Perturbed Orig Perturbed Orig Perturbed Accuracy Accuracy Accuracy
LLaVA-1.5-7B 95.3 43.5 97.0 57.4 62.4 56.4 46.3 45.2 64.5 64.3 63.4
   + VLMClassifier-1 imagewikiqa 15.1 0.0 15.6 0.0 61.5 61.5 47.8 47.5 61.1 36.2 35.9
   + VLMClassifier-2 imagewikiqa 15.6 0.0 15.1 0.0 61.8 61.2 47.4 47.8 61.8 35.8 36.0
Ours 98.6 98.4 99.3 98.9 81.8 81.0 51.5 51.0 67.8 73.1 64.6
LLaVA-1.5-13B 95.6 73.0 97.9 77.4 65.9 63.8 51.8 50.8 66.1 72.1 64.5
   + I-MoF mof 93.9 70.1 97.8 80.9 69.2 64.5 46.2 39.1 66.8 73.0 66.6
Ours 98.5 98.4 99.2 98.6 83.0 82.1 56.6 55.8 62.6 73.7 68.4
Achieving Pareto-Optimality Across Unimodal and Multimodal Tasks
As shown in Table 1, our method outperforms both VLMClassifier and I-MoF across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA (72.1%/64.5%) and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance—82.1% vs. 64.5% on OpenBookQA, and 55.8% vs. 39.1% on MMLU (LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
Table 2: Detailed multimodal reasoning accuracy (%) on multiple-choice VQA datasets across different ablation study settings. The best accuracy is marked in bold. Arrows (↑ / ↓) indicate relative changes compared to the Vanilla baseline of each model. Overall performance is computed as a weighted average across datasets, with weights proportional to each dataset’s test size.
Model Method ScienceQA-IMG MM-Bench-EN Seed-Bench-IMG VQA Overall
3B Multimodal Models
Qwen2.5-VL-3B qwen25vl Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 75.3 (↑19.5%) 82.0 (↑0.2%) 74.7 (↑3.3%) 76.1 (↑5.0%)
FFT with $D^{\text{AUG}}$ 73.6 (↑16.8%) 81.7 (↓0.1%) 75.3 (↑4.1%) 76.2 (↑5.1%)
   + KL 72.8 (↑15.6%) 80.9 (↓1.1%) 74.9 (↑3.6%) 75.9 (↑4.7%)
   + JS 73.4 (↑16.5%) 82.0 (↑0.2%) 75.2 (↑4.0%) 76.2 (↑5.1%)
   + RG 73.4 (↑16.5%) 81.8 (↑0.0%) 75.0 (↑3.7%) 76.0 (↑4.8%)
   + PGD 73.1 (↑16.0%) 81.7 (↓0.1%) 75.3 (↑4.1%) 76.1 (↑5.0%)
Ours 73.8 (↑17.1%) 81.5 (↓0.4%) 75.5 (↑4.4%) 76.4 (↑5.4%)
7B Multimodal Models
LLaVA-1.5-7B llava15 Vanilla 64.5 (100%) 64.3 (100%) 63.4 (100%) 63.8 (100%)
FFT with $D^{\text{VQA}}$ 61.6 (↓4.5%) 71.4 (↑11.0%) 62.4 (↓1.6%) 64.0 (↑0.3%)
FFT with $D^{\text{AUG}}$ 65.4 (↑1.4%) 71.1 (↑10.6%) 63.9 (↑0.8%) 65.6 (↑2.8%)
   + KL 65.9 (↑2.2%) 72.5 (↑12.8%) 64.5 (↑1.7%) 66.3 (↑3.9%)
   + JS 63.8 (↓1.1%) 73.5 (↑14.3%) 63.7 (↑0.5%) 65.6 (↑2.8%)
   + RG 66.3 (↑2.8%) 71.8 (↑11.7%) 63.7 (↑0.5%) 65.7 (↑3.0%)
   + PGD 66.7 (↑3.4%) 71.4 (↑11.0%) 63.6 (↑0.3%) 65.7 (↑3.0%)
Ours 67.8 (↑5.1%) 73.1 (↑13.7%) 64.6 (↑1.9%) 66.8 (↑4.7%)
InstructBlip-7B instructblip Vanilla 52.1 (100%) 65.5 (100%) 54.8 (100%) 56.4 (100%)
FFT with $D^{\text{VQA}}$ 61.8 (↑18.6%) 68.4 (↑4.4%) 57.6 (↑5.1%) 60.4 (↑7.1%)
FFT with $D^{\text{AUG}}$ 65.4 (↑25.5%) 71.1 (↑8.5%) 61.2 (↑11.7%) 63.9 (↑13.3%)
   + KL 65.9 (↑26.5%) 71.7 (↑9.5%) 60.9 (↑11.1%) 63.9 (↑13.3%)
   + JS 66.9 (↑28.4%) 72.3 (↑10.4%) 60.7 (↑10.8%) 64.1 (↑13.7%)
   + RG 62.8 (↑20.5%) 70.9 (↑8.2%) 62.5 (↑14.0%) 64.2 (↑13.8%)
   + PGD 66.0 (↑26.7%) 69.0 (↑5.3%) 61.8 (↑12.8%) 64.0 (↑13.5%)
Ours 64.0 (↑22.8%) 71.4 (↑9.0%) 63.0 (↑14.9%) 64.8 (↑14.9%)
13B Multimodal Models
LLaVA-1.5-13B llava15 Vanilla 65.8 (100%) 72.1 (100%) 64.5 (100%) 66.2 (100%)
FFT with $D^{\text{VQA}}$ 60.8 (↓7.6%) 73.6 (↑2.1%) 64.9 (↑0.6%) 65.8 (↓0.6%)
FFT with $D^{\text{AUG}}$ 63.5 (↓3.5%) 75.0 (↑4.0%) 65.7 (↑1.9%) 67.1 (↑1.4%)
   + KL 62.5 (↓5.0%) 74.6 (↑3.5%) 67.6 (↑4.8%) 68.0 (↑2.7%)
   + JS 65.8 (↑0.0%) 74.7 (↑3.6%) 67.6 (↑4.8%) 68.6 (↑3.6%)
   + RG 58.3 (↓11.4%) 73.5 (↑1.9%) 67.4 (↑4.5%) 66.9 (↑1.1%)
   + PGD 57.6 (↓12.5%) 74.2 (↑2.9%) 68.3 (↑5.9%) 67.5 (↑2.0%)
Ours 62.6 (↓4.9%) 73.7 (↑2.2%) 68.4 (↑6.0%) 68.4 (↑3.3%)
Ablation Studies
Table 3: Evaluation of Caltech-101 (Image-heavy) and MMLU (Text-heavy) across different ablation study settings. we report accuracy on the original setting (Orig) and the worst-performing perturbation (Perturbed).
Model Method Caltech-101 MMLU
Orig Perturbed Orig Perturbed
LLaVA-1.5-7B Vanilla 97.0 57.4 46.3 45.2
+$D^{\text{VQA}}$ 96.2 46.3 46.8 45.6
+$D^{\text{AUG}}$ 98.5 98.6 51.1 50.7
   +KL 98.6 98.7 51.1 50.7
   +PGD 98.7 98.5 50.6 50.3
Ours 99.3 99.0 51.5 51.0
InstructBLIP-7B Vanilla 90.3 17.5 35.3 35.2
+$D^{\text{VQA}}$ 92.1 23.1 40.9 40.2
+$D^{\text{AUG}}$ 99.0 56.1 50.0 49.7
   +JS 98.9 98.4 50.7 49.3
   +PGD 99.1 85.2 49.3 48.4
Ours 99.2 98.3 50.2 49.7
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise instead of adversarial perturbation), FFT+PGD (FFT with heuristic perturbations and adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table 2 presents overall VQA performance across different models and training strategies, while, Table 3 evaluates model robustness under unimodal settings by comparing accuracy on the original inputs versus the worst-performing perturbation cases. Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Table 4 reports results on unimodal tasks with all the perturbations.  Table 5 is full version of Table 3. Across all model families (Qwen2.5-VL, InstructBLIP, and LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +4.7% on LLaVA-1.5-7B, +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% on InstructBLIP-7B on Caltech101).  Figure 4 are radar chart visualizations. ( Table 4,  Table 5,  Figure 4 in  Appendix C)
We observe consistent improvements across both unimodal and multimodal tasks when moving from FFT w/ $D^{\text{VQA}}$ to FFT w/ $D^{\text{AUG}}$, highlighting the importance of incorporating modality-specific supervision and heuristic perturbations. For instance, on LLaVA-1.5-7B, both VQA accuracy (64.0% $\rightarrow$ 65.6%) and unimodal robustness (46.3% $\rightarrow$ 98.6% on Caltech101, 45.6% $\rightarrow$ 50.7% on MMLU) improve, indicating enhanced unimodal alignment. Building upon this, adding consistency regularization yields further gains by stabilizing model predictions under controlled perturbations on $X_{I}$ or $X_{T}$. Both KL and JS objectives lead to similar improvements, suggesting that the model equally benefits from all heuristic perturbations regardless of anchor choice. Finally, PGD-based adversarial perturbations outperform random Gaussian noise across almost all models, confirming that structured, learnable adversarial perturbation is more effective at weakening spurious modality shortcuts and encourages learning of more robust, task-relevant representations.
6 Conclusion
In this paper, we identify and formalize modality interference as a concrete manifestation of the broader cross-modality competency problem in Multimodal Large Language Models—namely, the inability to distinguish task-relevant from irrelevant modality signals. Through a designed perturbation-based causal evaluation experiment, we demonstrate that even state-of-the-art MLLMs systematically exhibit degraded performance under irrelevant but misleading inputs, revealing a fundamental vulnerability in their inference-time reasoning.
To mitigate this issue, we propose a robust fine-tuning strategy that combines modality-specific data augmentation, consistency regularization, and adversarial perturbation in the embedding space. These interventions explicitly constrain the model to produce stable outputs under spurious modality shifts, thereby reducing reliance on non-causal correlations and improving robustness. Extensive experiments across diverse architectures, scales, and task regimes confirm that our approach consistently improves both unimodal reasoning and multimodal generalization. Our findings suggest that enhancing cross-modality competency requires not only better alignment, but also causal safeguards against shortcut learning from irrelevant modalities.
Appendix A Appendix Summary
This appendix provides comprehensive supplementary materials and discussion to support the main findings of our paper on diagnosing and mitigating modality interference in MLLMs. We organize the appendix into several sections:
Finetuning StrategiesAppendix B): We elaborate on our design choice to freeze the Q-Former in InstructBLIP-based models. This decision is motivated by the need to retain strong visual representations while avoiding overfitting to perturbed or misleading multimodal inputs. ( Table 9 records the size for each dataset)
Detailed Experimental ResultsAppendix C): This section includes three key tables— Table 4,  Table 5 and  Table 6—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table 7 records the performance of different vanilla MLLMS under modality interference across modality-heavy datasets. We also include radar plots ( Figure 4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b [3] and InstructBlip-Vicuna-13b [2] and make further discussion on the selection of specific consistency loss.
Hyperparameter SettingsAppendix D): We present full training configurations used in our experiments, including optimization strategies, perturbation settings, and sampling ratios for different task types. This section enables reproducibility and highlights the computational efficiency of our proposed training scheme. We provide parameter analysis on iterations of PGD in  Figure 5.
Compute Resource DetailsAppendix E): We document hardware specifications, training durations, and resource costs for models of different scales. These details contextualize the feasibility of our approach in academic environments.
LimitationsAppendix F): We discuss the granularity of our current modality interference analysis, the selections of perturbations, and propose directions for more fine-grained future studies.
Broader ImpactsAppendix G): Finally, we reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
Together, these sections provide a complete view of our technical contributions, empirical findings, and responsible research considerations.
Table 4: Unimodal ability evaluation on image-heavy and text-heavy tasks under perturbation. Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas. The best accuracy is marked in bold.
Model Method Mini-ImageNet Caltech-101 OpenBookQA MMLU
Orig UF MD Orig UF MD RP RI FB FW RP RI FB FW
Qwen2.5-VL-3B [3] Vanilla 98.9 98.5 94.9 98.8 99.0 94.4 79.9 74.6 80.0 79.7 63.5 61.1 64.0 63.6
FFT with $D^{\text{VQA}}$ 98.8 98.5 95.3 98.8 98.1 94.3 80.7 74.3 80.1 80.2 63.0 61.7 63.0 63.3
FFT with $D^{\text{AUG}}$ 98.8 98.8 98.6 99.6 99.3 99.6 87.1 86.7 87.4 87.2 64.8 63.9 64.8 64.7
   + KL 98.9 98.7 99.1 99.6 99.5 99.7 87.1 86.2 86.7 86.8 66.0 65.5 65.9 65.9
   + JS 99.1 98.8 98.3 99.6 99.4 98.0 85.0 84.2 85.1 85.1 65.6 65.1 65.5 65.5
   + RG ($\sigma{=}0.05$) 99.0 99.1 98.9 99.5 99.5 99.2 86.4 86.9 86.9 87.2 64.6 64.3 64.6 64.8
   + PGD 99.3 99.3 99.1 99.5 99.4 99.5 86.6 85.8 86.8 86.6 65.3 64.0 65.4 65.3
Ours 99.3 99.2 99.2 99.7 99.7 99.5 86.7 86.4 86.6 86.6 64.8 64.5 65.0 65.1
Qwen2.5-VL-7B [3] Vanilla 99.3 99.3 96.3 99.1 98.9 97.2 85.9 77.5 85.8 86.0 69.3 63.7 68.9 68.9
FFT with $D^{\text{VQA}}$ 99.2 99.3 96.0 99.5 99.5 95.7 86.3 82.3 86.5 86.3 69.2 67.4 69.4 69.3
FFT with $D^{\text{AUG}}$ 99.6 99.5 99.4 99.7 99.7 99.5 90.2 90.2 90.3 90.3 70.4 69.9 70.4 70.3
   + KL 99.3 99.3 99.1 99.6 99.6 99.6 92.0 91.7 92.2 92.1 71.2 70.7 71.0 71.0
   + JS 99.5 99.4 99.4 99.7 99.3 99.6 91.6 92.1 92.2 92.1 71.5 69.9 71.5 71.6
   + RG ($\sigma{=}0.05$) 99.4 99.4 99.2 99.6 99.3 99.4 89.1 87.9 89.1 89.1 66.7 65.6 66.5 66.5
   + PGD 99.4 99.3 99.2 99.6 99.5 99.6 91.7 92.2 91.8 91.8 70.4 70.0 70.4 70.4
Ours 99.6 99.5 99.5 99.6 99.6 99.7 90.9 89.7 90.8 91.0 69.8 68.2 69.8 70.0
LLaVA-1.5-7B [13] Vanilla 95.3 93.4 43.5 97.0 95.9 57.4 62.4 56.4 62.5 63.4 46.3 45.2 45.9 45.8
FFT with $D^{\text{VQA}}$ 94.3 92.7 41.5 96.2 94.0 46.3 61.3 55.5 62.0 62.9 46.8 45.6 47.5 47.7
FFT with $D^{\text{AUG}}$ 98.2 98.2 98.1 98.5 98.6 99.0 78.6 77.2 78.7 78.4 51.1 50.7 51.1 51.3
   + KL 99.1 99.0 98.9 98.6 98.7 98.8 81.4 81.3 81.4 81.2 52.0 51.8 52.0 52.2
   + JS 98.7 98.8 99.0 99.1 99.0 99.2 81.6 81.6 81.7 81.5 51.6 51.8 52.4 52.4
   + RG ($\sigma{=}0.05$) 98.4 98.4 98.5 98.9 98.9 99.1 80.5 79.8 79.9 80.3 49.5 49.5 49.9 49.6
   + PGD 98.7 98.7 98.5 98.7 98.5 98.8 81.7 81.0 81.4 80.8 50.6 50.3 50.9 50.7
Ours 98.6 98.4 98.7 99.3 98.9 99.3 81.8 81.0 81.7 81.7 51.5 50.9 51.5 51.4
LLaVA-1.5-13B [13] Vanilla 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
FFT with $D^{\text{VQA}}$ 94.6 93.9 72.0 97.8 96.5 80.2 67.5 64.2 69.1 69.3 52.4 52.2 53.1 53.3
FFT with $D^{\text{AUG}}$ 98.1 96.8 98.4 96.7 96.9 97.0 81.0 78.7 81.1 81.3 52.1 51.7 51.8 51.6
   + KL 98.3 98.0 98.6 98.8 98.5 98.9 83.0 82.6 83.3 83.0 55.7 55.1 55.6 55.6
   + JS 98.3 98.1 98.0 98.7 98.4 98.7 83.1 81.5 83.1 83.1 56.7 56.2 56.6 56.5
   + RG ($\sigma{=}0.05$) 98.5 98.0s 98.1 98.9 98.5 98.9 83.5 82.5 83.1 82.8 55.4 55.3 55.7 55.5
   + PGD 98.7 98.2 98.6 99.0 98.6 99.0 82.2 82.6 82.6 82.8 55.6 55.4 55.6 55.5
Ours 98.5 98.4 98.7 99.2 98.6 99.2 83.0 82.1 82.7 83.1 56.7 55.8 56.7 56.7
InstructBlip-7B [2] Vinilla 92.0 87.1 13.6 90.3 90.2 17.5 50.8 46.2 50.9 50.7 35.3 35.8 35.2 35.7
FFT with $D^{\text{VQA}}$ 95.6 86.6 16.3 98.3 91.0 23.1 49.8 45.2 49.5 50.7 40.9 40.2 41.0 41.6
FFT with $D^{\text{AUG}}$ 98.5 98.0 38.2 99.0 98.7 56.1 75.0 74.9 74.8 75.8 50.0 49.7 50.0 50.0
   + KL 98.7 98.1 98.3 99.5 99.0 99.6 76.9 77.0 76.9 77.3 51.3 50.6 51.3 51.5
   + JS 98.5 97.7 98.5 98.9 98.4 98.9 78.0 76.6 77.7 78.0 50.7 50.1 50.7 50.8
   + RG ($\sigma{=}0.05$) 98.9 97.2 72.5 99.1 99.0 82.2 75.2 72.6 76.0 76.9 48.3 47.6 48.9 49.1
   + PGD 98.7 98.5 32.2 99.5 98.9 49.2 76.8 76.8 76.5 76.3 49.3 48.4 49.5 49.4
Ours 98.4 97.9 98.0 99.2 98.3 99.0 79.0 77.3 79.3 79.0 50.2 49.7 50.3 50.2
InstructBlip-13B [2] Vanilla 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
FFT with $D^{\text{VQA}}$ 95.6 85.8 8.0 97.0 87.5 11.6 58.6 55.4 59.7 60.6 43.7 42.8 43.6 44.1
FFT with $D^{\text{AUG}}$ 98.4 98.2 9.3 99.2 98.8 13.8 82.0 80.4 81.2 81.2 52.1 51.3 52.4 53.0
   + KL 98.5 98.3 98.7 99.1 99.2 99.5 82.5 81.4 82.1 82.9 53.4 52.5 53.4 53.4
   + JS 98.7 98.1 98.9 99.3 99.2 99.5 83.5 83.1 83.1 83.3 52.8 52.2 53.2 53.3
   + RG ($\sigma{=}0.05$) 98.4 97.8 87.0 99.4 99.3 94.4 80.0 76.6 79.6 80.4 50.9 50.0 51.4 51.8
   + PGD 98.6 98.0 80.9 98.7 98.6 99.1 79.8 79.0 80.9 80.9 51.3 50.7 51.4 52.4
Ours 98.7 97.9 98.0 98.7 98.7 98.8 83.2 81.2 83.8 83.0 52.2 51.6 52.3 53.4
Table 5: Evaluation of unimodal and multimodal tasks across different ablation study settings. For unimodal datasets, we report accuracy on the original setting (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on the original setting. The best accuracy is marked in bold.
Model Method Mini-ImageNet Caltech-101 OpenBookQA MMLU VQA Overall
Orig Perturbed Orig Perturbed Orig Perturbed Orig Perturbed Accuracy
3B Multimodal Models
Qwen2.5-VL-3B [3] Vanilla 98.9 94.9 98.8 94.4 79.9 74.6 63.5 61.1 72.5
FFT with $D^{\text{VQA}}$ 98.8 95.3 98.8 94.3 80.7 74.3 63.0 61.7 76.1
FFT with $D^{\text{AUG}}$ 98.8 98.6 99.6 99.3 87.1 86.2 64.8 63.9 76.2
   + KL 98.9 98.7 99.6 99.5 87.1 86.7 66.0 65.5 75.9
   + JS 99.1 98.3 99.6 98.0 85.0 84.2 65.6 65.1 76.2
   + RG 99.0 98.9 99.5 99.2 86.4 86.4 64.6 64.3 76.0
   + PGD 99.3 99.1 99.5 99.4 86.6 85.8 65.3 64 76.1
Ours 99.3 99.2 99.7 99.5 86.7 86.6 64.8 64.5 76.4
7B Multimodal Models
LLaVA-1.5-7B [13] Vanilla 95.3 43.5 97.0 57.4 62.4 56.4 46.3 45.2 63.8
FFT with $D^{\text{VQA}}$ 94.3 41.5 96.2 46.3 61.3 55.5 46.8 45.6 64.0
FFT with $D^{\text{AUG}}$ 98.2 98.1 98.5 98.6 78.6 77.2 51.1 50.7 65.6
   + KL 99.1 99.0 98.6 98.7 81.4 81.2 51.1 50.7 66.3
   + JS 98.7 98.8 99.1 99.0 81.6 81.5 52.0 51.8 65.6
   + RG 98.4 98.4 98.9 98.9 80.5 79.8 49.5 49.5 65.7
   + PGD 98.7 98.5 98.7 98.5 81.7 80.8 50.6 50.3 65.7
Ours 98.6 98.6 99.3 99.0 81.8 81.5 51.5 51.0 66.8
InstructBlip-7B [2] Vanilla 92.0 13.6 90.3 17.5 50.9 46.2 35.3 35.2 56.4
FFT with $D^{\text{VQA}}$ 95.6 16.3 98.3 23.1 49.8 45.2 40.9 40.2 60.4
FFT with $D^{\text{AUG}}$ 98.5 38.2 99.0 56.1 75.0 74.8 50.0 49.7 63.9
   + KL 98.7 98.1 99.5 99.0 76.9 76.9 51.3 50.6 63.9
   + JS 98.5 97.7 98.9 98.4 78.0 76.6 50.7 50.1 64.1
   + RG 98.9 72.5 99.1 82.2 76.0 72.6 48.3 47.6 64.2
   + PGD 98.7 72.2 99.5 85.2 76.8 76.3 49.3 48.4 64.0
Ours 98.4 98.0 99.2 98.3 79.0 78.3 50.2 49.7 64.8
Qwen2.5-VL-7B [2] Vanilla 99.3 96.3 99.1 97.2 85.9 77.5 69.3 63.7 80.3
FFT with $D^{\text{VQA}}$ 99.2 96.0 99.5 95.7 86.3 82.3 69.2 67.4 79.5
FFT with $D^{\text{AUG}}$ 99.6 99.4 99.7 99.5 90.2 90.2 70.4 69.9 79.9
   + KL 99.3 99.1 99.6 99.6 92.0 91.7 71.2 70.7 80.6
   + JS 99.5 99.4 99.7 99.3 91.6 92.1 71.5 69.9 80.3
   + RG 99.4 99.3 99.7 99.5 91.7 91.8 70.4 69.9 78.0
   + PGD 99.4 99.2 99.6 99.5 91.7 91.8 70.4 70.0 79.9
Ours 99.6 99.5 99.6 99.7 90.9 89.7 69.8 68.2 80.9
13B Multimodal Models
LLaVA-1.5-13B [13] Vanilla 95.6 73.0 97.9 77.4 65.9 63.8 51.8 50.8 66.2
FFT with $D^{\text{VQA}}$ 94.6 72.0 97.8 80.2 67.5 64.2 52.4 52.2 65.8
FFT with $D^{\text{AUG}}$ 98.1 96.8 96.7 96.9 81.0 78.7 52.1 51.6 67.1
   + KL 98.3 98.0 98.8 98.5 83.0 82.6 55.7 55.1 68.0
   + JS 98.3 98.1 98.7 98.4 83.1 81.5 56.7 56.2 68.6
   + RG 98.5 98.0 98.9 98.5 83.5 82.5 55.4 55.3 66.9
   + PGD 98.7 98.2 99.0 98.6 82.2 82.5 55.6 55.4 67.5
Ours 98.5 98.4 99.2 98.7 83.0 82.1 56.7 56.0 68.4
InstructBlip-13B [2] Vanilla 95.6 73.0 97.9 77.4 65.9 63.8 51.8 50.8 65.8
FFT with $D^{\text{VQA}}$ 95.6 8.0 97.0 11.6 58.6 55.4 43.7 42.8 59.4
FFT with $D^{\text{AUG}}$ 98.4 9.3 99.2 13.8 82.0 80.4 52.1 51.3 65.9
   + KL 98.5 98.3 99.1 99.2 82.5 81.4 53.4 52.5 66.2
   + JS 98.7 98.1 99.3 99.2 83.5 83.1 52.8 52.2 66.5
   + RG 98.4 97.8 99.4 99.3 80.0 76.6 50.9 50.0 66.3
   + PGD 98.6 80.9 98.7 98.6 79.8 79.0 51.3 50.7 66.4
Ours 98.7 97.9 98.7 98.7 83.2 81.2 52.2 51.6 66.5
Appendix B Finetuning Strategies
In our adaptation of InstructBLIP-Vicuna-7B, we choose to freeze the Q-Former and only fine-tune the language model and the projection layer. This decision is grounded in the nature of the Q-Former as a highly task-specific visual query encoder, originally pre-trained on VQA-style datasets where fine-grained and semantically aligned image-text pairs dominate.
However, in our setting, we deliberately introduce perturbations to the input modalities (e.g., injecting unrelated or misleading text/image content), which breaks the expected alignment structure. We observe that training the Q-Former under such noisy supervision leads to unstable representations and overfitting to spurious modality correlations. In contrast, freezing the Q-Former allows us to preserve its original strong visual grounding capabilities, while letting the downstream language model learn to filter or suppress misleading signals introduced during training.
This alternative tuning strategy enhances robustness under modality interference and aligns with our overall goal of improving cross-modal competency in MLLMs under perturbed conditions.
Appendix C Detailed Experimental Results
Please see  Table 4,  Table 5,  Figure 4,  Table 7,  Table 6 for more details.
Table 6: Detailed multimodal reasoning accuracy (%) on multiple-choice VQA datasets across different ablation study settings with extra models: Qwen2.5-vl-7B, Instructblip-Vicuna-13B. The best accuracy is marked in bold. Overall performance is computed as a weighted average across datasets, with weights proportional to each dataset’s test size.
Model Method ScienceQA-IMG MM-Bench-EN Seed-Bench-IMG VQA Overall
Qwen2.5-VL Models [3]
Qwen2.5-VL-3B [3] Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 75.3 (↑19.5%) 82.0 (↑0.2%) 74.7 (↑3.3%) 76.1 (↑5.0%)
FFT with $D^{\text{AUG}}$ 73.6 (↑16.8%) 81.7 (↓0.1%) 75.3 (↑4.1%) 76.2 (↑5.1%)
   + KL 72.8 (↑15.6%) 80.9 (↓1.1%) 74.9 (↑3.6%) 75.9 (↑4.7%)
   + JS 73.4 (↑16.5%) 82.0 (↑0.2%) 75.2 (↑4.0%) 76.2 (↑5.1%)
   + RG 73.4 (↑16.5%) 81.8 (↑0.0%) 75.0 (↑3.7%) 76.0 (↑4.8%)
   + PGD 73.1 (↑16.0%) 81.7 (↓0.1%) 75.3 (↑4.1%) 76.1 (↑5.0%)
Ours 73.8 (↑17.1%) 81.5 (↓0.4%) 75.5 (↑4.4%) 76.4 (↑5.4%)
Qwen2.5-VL-7B [3] Vanilla 85.1 (100%) 86.6 (100%) 77.0 (100%) 80.3 (100%)
FFT with $D^{\text{VQA}}$ 81.4 (↓4.3%) 86.4 (↓0.2%) 76.8 (↓0.3%) 79.5 (↓1.0%)
FFT with $D^{\text{AUG}}$ 83.2 (↓2.2%) 86.1 (↓0.6%) 77.0 (-0.0%) 79.9 (↓0.5%)
   + KL 86.0 (↑1.1%) 86.7 (↑0.1%) 77.2 (↑0.3%) 80.6 (↑0.4%)
   + JS 85.5 (↑0.5%) 85.9 (↓0.8%) 77.0 (-0.0%) 80.3 (-0.0%)
   + RG 81.6 (↓4.1%) 82.9 (↓4.2%) 75.5 (↓1.9%) 78.0 (↓2.8%)
   + PGD 85.2 (↑0.1%) 85.8 (↓0.9%) 76.6 (↓0.5%) 79.9 (↓0.5%)
Ours 83.9 (↓1.4%) 86.4 (↓0.2%) 78.3 (↑1.7%) 80.9 (↑0.7%)
Instructblip-Vicuna Models [2]
InstructBlip-7B [2] Vanilla 52.1 (100%) 65.5 (100%) 54.8 (100%) 56.4 (100%)
FFT with $D^{\text{VQA}}$ 61.8 (↑18.6%) 68.4 (↑4.4%) 57.6 (↑5.1%) 60.4 (↑7.1%)
FFT with $D^{\text{AUG}}$ 65.4 (↑25.5%) 71.1 (↑8.5%) 61.2 (↑11.7%) 63.9 (↑13.3%)
   + KL 65.9 (↑26.5%) 71.7 (↑9.5%) 60.9 (↑11.1%) 63.9 (↑13.3%)
   + JS 66.9 (↑28.4%) 72.3 (↑10.4%) 60.7 (↑10.8%) 64.1 (↑13.7%)
   + RG 62.8 (↑20.5%) 70.9 (↑8.2%) 62.5 (↑14.0%) 64.2 (↑13.8%)
   + PGD 66.0 (↑26.7%) 69.0 (↑5.3%) 61.8 (↑12.8%) 64.0 (↑13.5%)
Ours 64.0 (↑22.8%) 71.4 (↑9.0%) 63.0 (↑14.9%) 64.8 (↑14.9%)
InstructBlip-13B [2] Vanilla 65.8 (100%) 72.1 (100%) 63.8 (100%) 65.8 (100%)
FFT with $D^{\text{VQA}}$ 61.7 (↓6.2%) 68.5 (↓5.0%) 56.0 (↓12.2%) 59.4 (↓9.7%)
FFT with $D^{\text{AUG}}$ 66.6 (↑1.2%) 71.5 (↓0.8%) 64.0 (↑0.3%) 65.9 (↑0.2%)
   + KL 67.2 (↑2.2%) 72.6 (↑0.7%) 64.0 (↑0.3%) 66.2 (↑0.6%)
   + JS 67.9 (↑3.2%) 72.7 (↑0.8%) 64.2 (↑0.6%) 66.5 (↑1.1%)
   + RG 65.4 (↓0.6%) 73.5 (↑1.9%) 64.3 (↑0.7%) 66.3 (↑0.8%)
   + PGD 66.1 (↑0.5%) 74.0 (↑2.6%) 64.1 (↑0.5%) 66.4 (↑0.9%)
Ours 66.2 (↑0.6%) 73.2 (↑1.5%) 64.3 (↑0.7%) 66.5 (↑1.1%)
LLaVA1.5 Models [13]
LLaVA-1.5-7B [13] Vanilla 64.5 (100%) 64.3 (100%) 63.4 (100%) 63.8 (100%)
FFT with $D^{\text{VQA}}$ 61.6 (↓4.5%) 71.4 (↑11.0%) 62.4 (↓1.6%) 64.0 (↑0.3%)
FFT with $D^{\text{AUG}}$ 65.4 (↑1.4%) 71.1 (↑10.6%) 63.9 (↑0.8%) 65.6 (↑2.8%)
   + KL 65.9 (↑2.2%) 72.5 (↑12.8%) 64.5 (↑1.7%) 66.3 (↑3.9%)
   + JS 63.8 (↓1.1%) 73.5 (↑14.3%) 63.7 (↑0.5%) 65.6 (↑2.8%)
   + RG 66.3 (↑2.8%) 71.8 (↑11.7%) 63.7 (↑0.5%) 65.7 (↑3.0%)
   + PGD 66.7 (↑3.4%) 71.4 (↑11.0%) 63.6 (↑0.3%) 65.7 (↑3.0%)
Ours 67.8 (↑5.1%) 73.1 (↑13.7%) 64.6 (↑1.9%) 66.8 (↑4.7%)
LLaVA-1.5-13B [13] Vanilla 65.8 (100%) 72.1 (100%) 64.5 (100%) 66.2 (100%)
FFT with $D^{\text{VQA}}$ 60.8 (↓7.6%) 73.6 (↑2.1%) 64.9 (↑0.6%) 65.8 (↓0.6%)
FFT with $D^{\text{AUG}}$ 63.5 (↓3.5%) 75.0 (↑4.0%) 65.7 (↑1.9%) 67.1 (↑1.4%)
   + KL 62.5 (↓5.0%) 74.6 (↑3.5%) 67.6 (↑4.8%) 68.0 (↑2.7%)
   + JS 65.8 (↑0.0%) 74.7 (↑3.6%) 67.6 (↑4.8%) 68.6 (↑3.6%)
   + RG 58.3 (↓11.4%) 73.5 (↑1.9%) 67.4 (↑4.5%) 66.9 (↑1.1%)
   + PGD 57.6 (↓12.5%) 74.2 (↑2.9%) 68.3 (↑5.9%) 67.5 (↑2.0%)
Ours 62.6 (↓4.9%) 73.7 (↑2.2%) 68.4 (↑6.0%) 68.4 (↑3.3%)
Note. While Qwen2.5-VL was originally instruction-tuned with proprietary in-house data [3], our reproduced version uses only publicly available LLaVA instruction-tuning data. Even under this constraint and without access to VQA-specific tuning samples, our models achieve comparable or even better performance across all VQA datasets—highlighting the robustness and effectiveness of our proposed perturbation-consistent fine-tuning strategy.
Table 7: Performance (%) of Vinilla models under modality interference across four datasets. We show accuracy under clean (origin) and various perturbations: Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas.
Model Mini-ImageNet Caltech-101 Open-Book QA MMLU
Orig UF MD Orig UF MD RP RI FB FW RP RI FB FW
InternVL2-2B [38] 91.9 88.6 25.5 94.6 91.3 33.2 46.3 36.2 45.2 45.3 39.1 36.9 39.3 39.2
LLaVA-1.5-7B [13] 95.3 93.4 43.5 97.0 95.9 57.4 62.4 56.4 62.5 63.4 46.3 45.2 45.9 45.8
LLaVA-Next-7B [13] 93.4 90.0 28.5 97.0 93.4 31.6 54.6 52.9 55.9 55.5 45.9 45.0 45.9 45.8
LLaVA-1.5-13B [13] 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
InstructBlip-7B [2] 92.0 87.1 13.6 90.3 90.2 17.5 50.8 46.2 50.9 50.7 36.3 35.2 36.2 36.7
InstructBlip-13B [2] 93.0 81.6 50.9 94.1 82.7 51.0 44.7 39.4 45.6 46.1 40.1 37.4 40.5 42.9
QwenVL2-2B [39] 98.7 98.6 75.8 99.1 99.1 66.8 59.6 54.7 63.4 63.8 49.5 44.6 49.7 50.0
QwenVL2.5-3B [3] 98.9 98.5 94.9 98.8 99.0 94.4 79.9 74.6 80.0 79.7 63.5 61.1 64.0 63.6
QwenVL2-7B [39] 99.0 99.1 96.3 99.6 99.5 97.6 82.5 80.9 83.7 83.2 66.9 65.3 67.7 67.8
QwenVL2.5-7B [3] 99.3 99.3 96.5 99.3 99.4 96.8 86.2 80.9 86.3 86.4 69.5 67.6 69.1 69.1
C.1 Perturbation-based Evaluation Experiment Results
We conduct a controlled perturbation-based evaluation across various MLLMs, as shown in Table 7. Our results reveal that both vision and language tasks are vulnerable to cross-modal interference. In vision classification tasks, misleading textual descriptions (e.g., text contradicting image content) lead to severe performance drops. For example, InternVL2-2B and InstructBLIP-7B on Mini-ImageNet drop from 91.9% to 25.5% and from 92.0% to 13.6%, respectively. Conversely, for language tasks such as OpenBookQA and MMLU, irrelevant visual inputs—particularly semantically unrelated real images—also degrade performance. LLaVA-1.5-7B drops from 46.3% to 45.2% on MMLU, while InstructBLIP-13B sees over 5 points of degradation.
A consistent trend is that larger models exhibit greater robustness. Models like LLaVA-1.5-13B and QwenVL2.5-7B maintain high accuracy across all perturbation types—e.g., QwenVL2.5-7B sustains over 96% on Mini-ImageNet with misleading text and over 86% on OpenBookQA with irrelevant images—indicating improved modality disentanglement and reduced sensitivity to spurious correlations. Nonetheless, performance still degrades relative to clean inputs, highlighting that interference effects remain non-negligible even in stronger models.
C.2 Experiment Results on InstructBlip-Vicuna-13b and QwenVL-2.5-7b
To enable a more equitable comparison with existing multimodal models, we extend our method to two additional backbones: InstructBLIP-Vicuna-13B and QwenVL-2.5-7B. As shown in Table 6, our approach consistently improves performance across multiple VQA benchmarks, even under different instruction-tuning conditions, demonstrating its robustness and general applicability.
C.3 Discussion on KL&JS use for Consistency Regularization
Although both KL and JS divergence serve as effective objectives for consistency regularization, we find that JS consistently achieves slightly better results across most settings. Specifically, in both unimodal tasks (e.g., Mini-ImageNet, MMLU) and multimodal reasoning benchmarks (e.g., ScienceQA, SeedBench), JS-regularized models consistently outperform their KL counterparts by a small but observable margin. This trend holds across different model backbones and training configurations, including our final unified method (see “Ours” rows in Table 4 and 6). This suggests a marginal advantage of JS regularization in enhancing model robustness.
(a) Pretrained MLLMs
(b) LLaVA-1.5-7B
(c) LLaVA-1.5-13B
(d) InstructBLIP-7B
Figure 4: Task-wise robustness under perturbation. Each radar chart shows model accuracy (%) across Mini-ImageNet, Caltech-101 (image-heavy) and OpenBookQA, MMLU (text-heavy) under various perturbations. (a) uses raw accuracy of different pretrained MLLMs directly. (b–d) are normalized relative accuracy of each MLLMs. (We normalize each absolute accuracy into relative accuracy, which refers to absolute tested accuracy / accuracy of vanilla MLLMs in origin setting without perturbation.)
Appendix D Hyper-Parameter Setting and Training details
Please see  Table 9 for more details.
Table 8: Hyperparameter Settings
Category Setting
Model and Training Strategy
Base Model LLaVA-1.5-7B
Finetune Type Full
Adversarial Type PGD
Step size $\alpha$ 0.1
Epsilon-hall $\epsilon$ 0.001
PGD Steps $T$ 2
Consistency Regularization Type JS
Loss Weight $\lambda_{\text{consistency}}$ 0.01
Temperature $\tau$ 1
Optimization
Epochs 1
Batch Size per GPU $|\mathcal{B}|$ 8
Img/Text Ratio 0.25/0.25
Learning rate $2\times 10^{-5}$
Table 9: Dataset Statistics
Dataset Train Test
Mini-ImageNet 4935 2000
Caltech-101 8124 1020
OpenBookQA 4957 1000
MMLU 7M 5469
LLaVA-Instruct 624610 -
TextCaps 109765 -
MMBench-EN - 4377
ScienceQA-IMG - 4114
SeedBench-IMG - 14243
D.1 Parameter Analysis with PGD iterations
To investigate the effect of adversarial strength on model performance, we vary the number of PGD training iterations from 1 to 5 and evaluate the resulting VQA accuracy. As shown in  Figure 5, both LLaVA-1.5-7B and LLaVA-1.5-13B models benefit from PGD-based consistency training, with performance peaking at 2-step PGD (66.82% and 68.37%, respectively). Notably, excessive iterations (e.g., 4 or 5 steps) may lead to slight degradation, especially in larger models, likely due to over-perturbation and optimization difficulty.
These findings suggest that a moderate PGD setting (2 steps with $\epsilon$=1e-3 and $\alpha$=0.1 in LLaVA-1.5-7b, $\epsilon$=1e-4 and $\alpha$=0.1 in LLaVA-1.5-13b) offers an optimal balance between robustness and training stability, and that model size influences sensitivity to adversarial signal strength.
Figure 5: Comparison of VQA performance across PGD iterations for different model sizes.
Appendix E Experiments compute resources
All experiments were conducted on 4×A6000 GPUs using DeepSpeed ZeRO-3 [40] with offloading. For 3B and 7B models, standard full fine-tuning takes approximately 8 and 10 hours, respectively. Adding KL regularization increases 7B training time to 15 hours, while introducing PGD-based perturbations further extends it to 18 hours (18 hours for 1 iteration, adding one iteration will add 2 hours). For larger-scale experiments with LLaVA-1.5-13B, we use 8×H100 GPUs with DeepSpeed ZeRO-2, where full fine-tuning completes in approximately 4 hours. We estimate the total compute cost to be within the typical academic scale.
Appendix F Limitations
Our analysis of modality interference is conducted from a coarse-grained perspective, primarily categorizing tasks into image-heavy and text-heavy types. A more fine-grained investigation—such as dynamic attention tracking—could provide deeper insights into how MLLMs rely on or ignore specific modalities during reasoning.
Moreover, while our perturbation strategies (e.g., unrelated facts, misleading descriptions, irrelevant images) effectively reveal failure modes of current MLLMs, they remain heuristic and task-specific. Designing perturbations is, by nature, an open-ended process—one can always propose new forms of misleading inputs. Thus, an ultimate goal is to develop perturbation-agnostic methods that improve robustness without requiring exhaustive enumeration of possible attacks.
While our use of PGD-based adversarial training represents a strong and generalizable perturbation strategy, it still operates within a defined input space (e.g., embedding-level noise bounded by $L_{\infty}$ norms). Hence, PGD-based adversarial perturbation should be viewed as a practical but partial solution rather than a comprehensive defense. Developing mechanisms that generalize across both semantic and modality perturbations remains an open and challenging direction.
Appendix G Broader Impacts
This work investigates the limitations of current multimodal large language models in reasoning across modalities and proposes methods to mitigate modality interference—a concrete failure case of cross-modality competency. By improving the robustness and alignment behavior of MLLMs, our approach may benefit a variety of downstream applications that rely on accurate visual-linguistic understanding, including education, accessibility tools (e.g., visual question answering for blind users), and scientific multimodal reasoning tasks.
Our findings also highlight the hidden risks of over-relying on irrelevant modality signals, which can degrade performance or lead to misleading predictions. Making such failure modes measurable and diagnosable can support safer deployment and more transparent evaluation of MLLMs in practice.
On the other hand, techniques such as adversarial perturbation (e.g., PGD-based finetuning) may be dual-use. While our implementation uses perturbations to improve model alignment, similar strategies could be misused to manipulate model behavior. To reduce such risks, we restrict all experiments to open-source academic models and do not include harmful or sensitive content in training or evaluation. We encourage future work to further assess modality interference in safety-critical contexts and to investigate alignment-aware perturbation techniques with explicit safety constraints.